# Realigning Incentives to Build Better Software: a Holistic Approach to Vendor Accountability

### v1

@@ 3.3 Software Liability Insurance
In terms of limits, Cowbell offers Tech E&O Excess coverage up to \$5M in limits for mid-sized companies with revenue up to \$1B,666see https://cowbell.insure/news-events/pr/cowbell-us-upmarket, and At-Bay provides coverage with aggregate limits of \$10M, for businesses with revenue up to \$5B777see https://www.at-bay.com/press_releases/expands-cyber-tech-coverage/, while Resilience offers limits up to \$10M for firms with revenue from \$300M to \$10B.888see https://www.insurancebusinessmag.com/us/news/cyber/resilience-adds-tech-eando-cover-to-cyber-proposition-480136.aspx
@@ 3.4 Regulatory Body
Secondly, we advocate for the establishment of an office that collects consumer complaints against software defects and data breaches stemming from software defects, just as the FTC’s Consumer Sentinel and FBI’s Internet Crime Complaint Center (IC3) collect consumer reports regarding identity theft and fraud reports. These reports would allow consumers to report the nature of the harm they suffered due to faulty software, such as lost work, lost income, physical and mental health issues, etc. Indeed, attempting to measure the actual costs of software defects will greatly help inform our collective understanding of the impact of harms.
@@
We end this section by briefly commenting on how our framework differs from the current system in the EU. The EU regulation may be summed up as a combination of (1) mandatory product and organizational audit and certification and (2) strict product liability (without upper limit). The system we advocate, on the other hand, consists of (i) product/process audits and (ii) the floor-ceiling hybrid approach to liability, which is a weaker version of (2). We also generally believe these should be voluntary, with the possible exception of the floor. An established floor could be made mandatory, meaning no product without minimal floor audit can enter the market, while the ceiling is voluntary (resulting in liability waiver or capped liability). We believe the gap between our liability model with non-mandatory audit and the EU system can be effectively filled by litigation and insurance, as detailed above.
@@ 4.3 Root cause analysis
After the crash, the machine either entered a boot loop or booted into recovery mode; either way, manual intervention was needed to restore proper operation. Given the huge number of host machines affected at the same time, the impact was global. Note that the so-called incremental rollout (or canary release999referring to the use of canaries in coal mines as sentinels for toxic gases) is an existing software best practice [42]; by using such a method, the vendor could observe the immediate impact of the newly released update on a smaller set of machines, before gradually pushing the update to more hosts. Such graceful deployment is (would have been) especially beneficial for critical, kernel-level software.
@@ 4.4 Legal actions
There has been significant backlash from customers, shareholders, and regulators [^10] https://www.securityweek.com/crowdstrike-faces-lawsuits-from-customers-investors/. The most prominent attached lawsuit is certainly the one filed by Delta Airlines, which is estimated to have suffered a loss of USD 500 million (not including reputation issues and potential TSA fines) and has had its hands full with more than 176,000176000176,000176 , 000 refund/reimbursement requests. The airline has hired a prominent attorney to pursue potential damages from CrowdStrike and, interestingly, Microsoft. Although Delta is believed to be seeking to recoup all its losses, CrowdStrike moved to dismiss the lawsuit, citing a contractual limit on its own liability and a cap on damages [^11] https://www.cnbc.com/2024/12/17/crowdstrike-moves-to-dismiss-delta-suit-citing-contract-terms.html. Furthermore, CrowdStrike filed a countersuit, blaming Delta’s own flawed internal processes for the majority of flight cancellations and refusing their technical assistance during the outage. (Note that other airlines were indeed more effective in handling the outage.) Microsoft made similar comments on grounds of Delta refusing technical assistance and having an outdated IT infrastructure.
@@
Lastly, CrowdStrike’s CEO was called to testify in front of the US Congress, and another senior executive apologized multiple times [^13] https://www.theguardian.com/technology/2024/sep/24/crowdstrike-outage-microsoft-apology for letting their customers down, and took “full responsibility” for the crashes. Although CrowdStrike conducted an honest forensic root cause analysis[41] with public results and vowed to improve their software-related processes including implementing incremental rollout, evidently, it does not admit legal responsibility, i.e., liability.
@@ 4.5 What if …
One could hypothesize that CrowdStrike’s issues stem from insufficient technical expertise; although human error is always a factor, it is far more likely that their simplified software-related procedures were the results of business decisions. They focused mainly on quickly releasing new product features, keeping the customers happy and their own testing costs low. Simply put, there were no sufficient incentives present to make thorough testing profitable; a textbook incentive issue in software security [^14] That CrowdStrike is a security vendor makes the situation a bit more striking. Should our proposed policy framework be adopted, proper incentives for paying attention to (security) testing would emerge: only by passing the product-based floor audit could they market their product; furthermore, successfully passing the voluntary process-based ceiling would result in waived/capped liability, while skipping or failing it would bring with itself full legal responsibility.
@@
Our policy framework (assuming proper regulations are passed) provides a much clearer legal situation compared to today’s US landscape [17]. With the uncertainty gone from legal outcomes, the financial resources of affected stakeholders could actually go into making safer software. Moreover, CrowdStrike’s own forensic investigation and pledge to engage independent third parties to conduct a further review of their quality control and software release processes essentially constitute a costly post-incident audit. It is reasonable to assume that an actual pre-incident audit would have been socially optimal as all impacted stakeholders (CrowdStrike and its shareholders, Microsoft, Delta and other customers, flight passengers and other clients of customers, etc.) could have been better off.
Lastly, although CrowdStrike lost 44% of its value over the 2 weeks after the incident, its share price has almost returned to pre-incident level in the next 2 months [^15] https://www.investing.com/news/analyst-ratings/crowdstrike-stock-outlook-revised-upward-sector-perform-rating-held-amid-nearterm-risks-93CH-3744295. Somewhat surprisingly, the company reported reasonably good 2024 Q3 numbers with an even better outlook for Q4161616https://ir.crowdstrike.com/news-releases/news-release-details/crowdstrike-reports-third-quarter-fiscal-year-2025-financial/. They attributed the financial bounce-back to their superior technology, customer commitment packages, and regained reputation. However, even in a market with several competitors, enterprise security software is a sticky product; there is a considerable lock-in effect, especially for larger customers with complex IT infrastructures. This has significantly contributed to a “monoculture” in cybersecurity products; its persistence will only make the next event more costly, more disruptive, and more wide-spread. Even without knowing the future fortune of the company, such a bounce-back from a major blunder on the back of imperfect competition carries a questionable message. The proposed policy framework could steer the ecosystem in more desirable directions.
@@ 5.2 AI (Software) Liability
The enormous risk AI presents is most recently exemplified in a case [43] involving a 14-year-old in Florida who committed suicide after allegedly being encouraged by an AI chatbot. This type of harm notwithstanding, we argue that in order to define, recognize, regulate, and mitigate AI risks, one must first do the same for software risks more generally. Without this starting point, we would quickly find ourselves in a situation where we must debate what differentiates AI software from non-AI software and which is governed by what type of regulation. Furthermore, even if there is agreement on the end AI product being distinctly different from other types of software, one must also consider the long chain of software components and mechanisms that went into producing the end product: data acquisition and curation, filtering, basic mathematical algorithms, post-processing, etc. Many of these mechanisms (as implemented in software) are not at all AI-specific – one finds them in millions of other clearly non-AI systems. Where should one draw the line?
@@
In short, we think it would be helpful and productive for the many stakeholders to first take the high-level view that AI is no more or less than a particular type of software. We do acknowledge the need in the software liability ecosystem to further carve out a space to address AI liability more specifically in future work. The set of corresponding EU software/AI regulatory measures (CRA, PLD, AI Act, and the draft AILD) aim to achieve just that.
On a related note, AI systems deserve special attention in terms of the new risks they pose. In fact, the AI Act [44] classifies AI systems into risk categories and regulates them accordingly. Stakeholders involved (suppliers, importers, deployers, etc,) must adhere to specific requirements, with special emphasis on the high-risk category. This regulation aims to prevent AI-related incidents, while the AILD [16] assigns liability in this context. On the contrary, President Trump has just rescinded the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence (EO 14110 [45]) and signed a new Executive Order on Removing Barriers to American Leadership in Artificial Intelligence (EO 14179 [46]). This move signals a move away from safety and towards minimal barriers to foster innovation and U.S. leadership in AI. One might wonder whether an alternative approach, a type of Hippocratic oath for AI developers [47], could be successful in the face of minimal regulations and ineffective litigation against big tech.
@@ 5.3 Limitations
While we believe the framework we have outlined is sound and can effectively help re-align crooked incentives, we also acknowledge that “the devil is in the details”: much remains to be worked out for each of the components in the ecosystem. Converging on a floor-safe harbor legal regime will require more specific definitions of the floor and the safe harbor; scholars will need to agree on what constitutes minimum requirement and what constitutes reasonable best effort. How to perform software audit is another major open technical area: is it possible to establish a set of audit standards, how to ensure this does not become yet another check-list, when a certain software has been certified to enter the clearinghouse, is it supposed to be safe for all use cases? In this sense, we view this paper as a call to arms: this effort will require the engagement of a large research community, and the time to act is now.
@@ 6 Conclusion
In a very recent and positive development, the outgoing U.S. administration passed Executive Order 14144 [48], which builds on EO 14028 [19] to strengthen cybersecurity and promote (software) innovation at the same time. Among others, the new EO sets out mandatory software security requirements for suppliers of the federal government: software providers must submit machine-readable attestations of secure development practices, along with validation artifacts, to the Cybersecurity and Infrastructure Security Agency (CISA). If such stringent requirements can spill over to the non-government software market, they have the potential to catalyze the proposed incentive re-alignment. Note that this EO has not been rescinded, seemingly affirming that software security is a bi-partisan issue.
@@ A.1 Software liability
Ex-post liability is built on waiting for an adverse event to occur, then assigning liability to the injurer [49]. The legal interpretation of liability already appeared in Roman Law, but the modern framework of product liability first emerged in the early twentieth century [50] in relation to a faulty automobile wheel causing injuries to the driver. The ruling cited the reasonable basis for the automobile manufacturer to know product risks and to make extra effort to ensure the safety of anyone coming into contact with the car, even if the wheel was produced by another company. The next milestone case in 1932 [51] established duty of care with respect to a person who consumed spoiled ginger beer produced by a bottler and contracted severe gastroenteritis. The bottler had to pay restitution owing to his negligence. Finally, standard of care was put forward in 1947  [52] in a complicated multi-party liability case among the US government, a railroad company, and two boating companies, where the mooring lines of a barge carrying flour belonging to the government were disconnected by another barge, causing damage to other ships and eventually sinking. Liability was shared among the companies; specifically, the boating company operating the barge that cut the lines failed to look after the safety and security of other barges contacted. Interestingly, the judge based his ruling on inadequate risk management: the burden of precautions taken by the company was smaller than the likelihood, and the impact of the incident multiplied together.
@@
Although there were several liability cases against software companies in the last decades, companies often escaped fines using nuances of contract law and limitations of liability and disclaimers in end-user license agreements (EULA) [17]. In fact, the flexible interpretation of acceptable risk taken by the user of a software product made this a gray area. This has changed dramatically towards the end of the last decade with the emergence of the European General Data Protection Regulation (GDPR) and its penalty system for organizations mismanaging the personal data of their users. The GDPR has defined both the duty and standard of care, data protection agencies have collected billions of Euros in fines, and the cost-benefit assessment of cybersecurity-related investments has changed for the better, at least in terms of the safekeeping of personal data. Concurrently, governments worldwide have become increasingly worried about the changing landscape of cyberattacks which showed a shift from accessing/exposing user data towards damaging and controlling critical infrastructure integrated across governmental institutions and private organizations. The 2017 NotPetya ransomware [53] and the 2021 SolarWinds attacks [54] stand out as having far-reaching implications for national security, letting malicious attackers breach and potentially take over control of systems of critical importance (governmental organizations, financial institutes, logistics companies, etc.) in the Ukraine, in the United States, and across the globe. These incidents made governments prioritize the cyber-hardening of their national infrastructure and, specifically, focus on software supply chain security. Accordingly, both the United States (via the 2021 Executive Order on Improving the Nation’s Cybersecurity [19]) and the European Union (via the 2022 Cyber Resilience Act [55]) acted swiftly and put forward regulations mandating secure software development, standardized product cybersecurity compliance, and a secure software supply chain [56], and also reducing the information asymmetry pertaining to non-expert end-users of software products. Software liability also factors in the end-user’s technical (lacking expertise to assess product security) and legal inability (EULA prohibiting the inspection of software code) to evaluate software products they use. In fact, the EU has enacted a new Product Liability Directive [15] that came into force on 9th December 2024, with the intent to explicitly designate software as a “product”; this, of course, comes with strict liability clauses. The recent CrowdStrike incident has shown the world that this just might be an effective approach for managing such scenarios (see Section 4 for a detailed case study).
@@
Another intriguing aspect of software liability comes from cyber-insurance. The devastating cybersecurity incidents and the ongoing multi-modal armed conflicts affecting critical infrastructure raised alarms in the insurance industry regarding the insurability of cyber risks [57], potentially inhibiting the market that experienced real growth in recent years after decades of underdevelopment. In fact, insurers, re-insurers, and academics and cybersecurity experts are trying to establish baseline scenarios in the industrial control system domain [58]. Indeed, if the cyber risk is systemic and catastrophic losses are expected in the critical infrastructure sector, the insurance industry will lose interest in underwriting new policies. In that case, governments would shoulder the responsibility to be insurers of last resort, similar to instances of frequent natural disasters [59]. Supposing this happens, the increased scrutiny for software product security laid down in recent regulations is even more sensible [60]. Even so, a potential insurer role would also be a financial and administrative burden for the government, especially taking into account tangled software supply chains. On the other hand, supply chain security solutions and best practices are available; if a company decides to invest money and effort, a much-improved level of security is within reach. Recognizing this, the current United States National Cybersecurity Strategy [61], released in April 2023, proposed a liability waiver mechanism designed to reward software companies willing to undergo and pass a government-mandated product security audit. (Such audits have been studied extensively; we refer the reader to [62] for an overview.) Such a waiver mechanism could just be the missing financial incentive for software companies to change their usual ways regarding secure products.

### v2

@@ 3.3 Software Liability Insurance
In terms of limits, Cowbell offers Tech E&O Excess coverage up to \$5M in limits for mid-sized companies with revenue up to \$1B,777see https://cowbell.insure/news-events/pr/cowbell-us-upmarket, and At-Bay provides coverage with aggregate limits of \$10M, for businesses with revenue up to \$5B888see https://www.at-bay.com/press_releases/expands-cyber-tech-coverage/, while Resilience offers limits up to \$10M for firms with revenue from \$300M to \$10B.999see https://www.insurancebusinessmag.com/us/news/cyber/resilience-adds-tech-eando-cover-to-cyber-proposition-480136.aspx
@@ 3.4 Regulatory Body
Secondly, we advocate for the establishment of an office that collects consumer complaints against software defects and data breaches caused by software defects, just as the FTC’s Consumer Sentinel and the FBI’s Internet Crime Complaint Center (IC3) collect consumer reports regarding identity theft and fraud reports. These reports would allow consumers to report the nature of the harm they suffered due to faulty software, such as lost work, lost income, physical and mental health issues, etc. Indeed, attempting to measure the actual costs of software defects will greatly help inform our collective understanding of the impact of harms.
@@
We end this section by briefly commenting on how our framework differs from the current system in the EU. The EU regulation may be summed up as a combination of (1) mandatory product and organizational audit and certification, and (2) strict product liability (without upper limit). The system we advocate, on the other hand, consists of (i) product/process audits and (ii) the floor-ceiling hybrid approach to liability, which is a weaker version of (2). We also generally believe these should be voluntary, with the possible exception of the floor. An established floor could be made mandatory, meaning no product without a minimal floor audit can enter the market, while the ceiling is voluntary (resulting in liability waiver or capped liability). We believe the gap between our liability model with non-mandatory audit and the EU system can be effectively filled by litigation and insurance, as detailed above.
@@ 4.3 Root cause analysis
After the crash, the machine either entered a boot loop or booted into recovery mode; either way, manual intervention was needed to restore proper operation. Given the huge number of host machines affected at the same time, the impact was global. Note that the so-called incremental rollout (or canary release101010referring to the use of canaries in coal mines as sentinels for toxic gases) is an existing software best practice [47]; by using such a method, the vendor could observe the immediate impact of the newly released update on a smaller set of machines, before gradually pushing the update to more hosts. Such graceful deployment is (would have been) especially beneficial for critical, kernel-level software.
@@ 4.4 Legal actions
There has been significant backlash from customers, shareholders, and regulators [^11] https://www.securityweek.com/crowdstrike-faces-lawsuits-from-customers-investors/. The most prominent attached lawsuit is certainly the one filed by Delta Airlines, which is estimated to have suffered a loss of USD 500 million (not including reputation issues and potential TSA fines) and has had its hands full with more than 176,000176000176,000176 , 000 refund/reimbursement requests. The airline has hired a prominent attorney to pursue potential damages from CrowdStrike and, interestingly, Microsoft. Although Delta is believed to be seeking to recoup all its losses, CrowdStrike moved to dismiss the lawsuit, citing a contractual limit on its own liability and a cap on damages [^12] https://www.cnbc.com/2024/12/17/crowdstrike-moves-to-dismiss-delta-suit-citing-contract-terms.html. Furthermore, CrowdStrike filed a countersuit, blaming Delta’s own flawed internal processes for the majority of flight cancellations and refusing their technical assistance during the outage. (Note that other airlines were indeed more effective in handling the outage.) Microsoft made similar comments on the grounds of Delta refusing technical assistance and having an outdated IT infrastructure.
@@
Lastly, CrowdStrike’s CEO was called to testify in front of the US Congress, and another senior executive apologized multiple times [^14] https://www.theguardian.com/technology/2024/sep/24/crowdstrike-outage-microsoft-apology for letting their customers down, and took “full responsibility” for the crashes. Although CrowdStrike conducted an honest forensic root cause analysis[46] with public results and vowed to improve their software-related processes, including implementing incremental rollout, evidently, it does not admit legal responsibility, i.e., liability.
@@ 4.5 What if …
One could hypothesize that CrowdStrike’s issues stem from insufficient technical expertise; although human error is always a factor, it is far more likely that their simplified software-related procedures were the result of business decisions. They focused mainly on quickly releasing new product features, keeping the customers happy, and their own testing costs low. Simply put, there were no sufficient incentives present to make thorough testing profitable; a textbook incentive issue in software security [^15] That CrowdStrike is a security vendor makes the situation a bit more striking. Should our proposed policy framework be adopted, proper incentives for paying attention to (security) testing would emerge: only by passing the product-based floor audit could they market their product; furthermore, successfully passing the voluntary process-based ceiling would result in waived/capped liability, while skipping or failing it would bring with itself full legal responsibility.
@@
Our policy framework (assuming proper regulations are passed) provides a much clearer legal situation compared to today’s US landscape [20]. With the uncertainty gone from legal outcomes, the financial resources of affected stakeholders could actually go into making safer software. Moreover, CrowdStrike’s own forensic investigation and pledge to engage independent third parties to conduct a further review of their quality control and software release processes essentially constitute a costly post-incident audit. It is reasonable to assume that an actual pre-incident audit would have been socially optimal, as all impacted stakeholders (CrowdStrike and its shareholders, Microsoft, Delta, and other customers, flight passengers, and other clients of customers, etc.) could have been better off.
Lastly, although CrowdStrike lost 44% of its value over the 2 weeks after the incident, its share price has almost returned to pre-incident level in the next 2 months [^16] https://www.investing.com/news/analyst-ratings/crowdstrike-stock-outlook-revised-upward-sector-perform-rating-held-amid-nearterm-risks-93CH-3744295. Somewhat surprisingly, the company reported reasonably good 2024 Q3 numbers with an even better outlook for Q4171717https://ir.crowdstrike.com/news-releases/news-release-details/crowdstrike-reports-third-quarter-fiscal-year-2025-financial/. They attributed the financial bounce-back to their superior technology, customer commitment packages, and regained reputation. However, even in a market with several competitors, enterprise security software is a sticky product; there is a considerable lock-in effect, especially for larger customers with complex IT infrastructures. This has significantly contributed to a “monoculture” in cybersecurity products; its persistence will only make the next event more costly, more disruptive, and more widespread. Even without knowing the future fortune of the company, such a bounce-back from a major blunder on the back of imperfect competition carries a questionable message. The proposed policy framework could steer the ecosystem in more desirable directions.
@@ 5.2 AI (Software) Liability
The enormous risk AI presents is most recently exemplified in a case [48] involving a 14-year-old in Florida who committed suicide after allegedly being encouraged by an AI chatbot. This type of harm notwithstanding, we argue that in order to define, recognize, regulate, and mitigate AI risks, one must first do the same for software risks more generally. Without this starting point, we would quickly find ourselves in a situation where we must debate what differentiates AI software from non-AI software and which is governed by what type of regulation. Furthermore, even if there is agreement on the end AI product being distinctly different from other types of software, one must also consider the long chain of software components and mechanisms that go into producing the end product: data acquisition and curation, filtering, basic mathematical algorithms, post-processing, etc. Many of these mechanisms (as implemented in software) are not at all AI-specific – one finds them in millions of other clearly non-AI systems. Where should one draw the line?
@@
In short, we think it would be helpful and productive for the many stakeholders to first take the high-level view that AI is a particular type of software. We do acknowledge the need in the software liability ecosystem to further carve out a space to address AI liability more specifically. AI systems deserve special attention because they indeed pose new types of risks. For instance, the random and erratic ways LLMs can behave (e.g., in response to prompt engineering) bring additional uncertainty. Another thought-provoking issue is the fact that such systems are only partially “patchable” in the conventional sense; new model versions and their dependence and interaction with previous iterations can exhibit deeper and more complex patterns beyond the capabilities of regular software updates. The issues above suggest that AI risks can ultimately differ in profound ways from regular software.
The set of interconnected EU software/AI regulatory measures (CRA, PLD, AI Act, and the now-revoked draft AILD) aims to achieve AI-specific software liability. In particular, the AI Act [49] classifies AI systems into risk categories and regulates them accordingly. Stakeholders involved (suppliers, importers, deployers, etc.) must adhere to specific requirements, with special emphasis on the high-risk category. This regulation aims to prevent AI-related incidents, while the draft AILD [19] assigned liability in this context. In the U.S., President Trump has rescinded the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence (EO 14110 [50]) and signed a new Executive Order on Removing Barriers to American Leadership in Artificial Intelligence (EO 14179 [51]). This move signals a move away from safety and towards minimal barriers to foster innovation and U.S. leadership in AI.
Interestingly, in the spring of 2025, the EU withdrew the draft AILD, citing the exact “fostering innovation” angle. Meanwhile, the California Senate has been discussing SB-813, a bill that proposes a landmark safe harbor for AI developers. This constitutes a (partial) waiver against civil liability claims should the AI developer undergo successful certification conducted by state-approved but independent Multistakeholder Regulatory Organizations (MROs). With international AI legislation seemingly in flux, one might wonder whether an alternative approach, a type of Hippocratic oath for AI developers [52], could be successful in the face of minimal regulations and ineffective litigation against big tech.
@@ 5.3 Limitations
While we believe the framework we have outlined is sound and can effectively help realign misaligned incentives, we also acknowledge that “the devil is in the details”: much remains to be worked out for each of the components in the ecosystem. In particular:
  • Converging on a floor-safe harbor legal regime will require more specific definitions of the floor and the safe harbor; scholars will need to agree on what constitutes a minimum requirement and what constitutes reasonable best effort.
  • How to perform a software audit is another major open technical area: is it possible to establish a set of audit standards, how to ensure this does not become yet another checklist, when a certain software has been certified to enter the clearinghouse, is it supposed to be safe for all use cases?
  • How to reconcile the discrete (one-time) nature of audit and the far more fluid nature of software development: can a vendor submit one version for audit but then release onto the market a different version (e.g., by turning off privacy-enhancing features)?
  • How to balance the accuracy and efficacy of an audit with the privacy/proprietary needs of the vendor?
In this sense, we view this paper as a call to arms: this effort will require the engagement of a large research community, and the time to act is now.
@@ 6 Conclusion
In a very recent and positive development, the outgoing U.S. administration passed Executive Order 14144 [53], which builds on EO 14028 [22] to strengthen cybersecurity and promote (software) innovation at the same time. Among others, the new EO sets out mandatory software security requirements for suppliers of the federal government: software providers must submit machine-readable attestations of secure development practices, along with validation artifacts, to the Cybersecurity and Infrastructure Security Agency (CISA). If such stringent requirements can spill over to the non-government software market, they have the potential to catalyze the proposed incentive realignment. Note that this EO has not been rescinded, seemingly affirming that software security is a bipartisan issue.
Acknowledgments
Gergely Biczók was supported by the Fulbright Visiting Scholar program and Project no. 138903 of the Ministry of Innovation and Technology, Hungary, from the NRDI Fund, financed under the FK_21 funding scheme. Mingyan Liu was supported by the US National Science Foundation (NSF) under grant IIS-2112471.
@@ A.1 Software liability
Ex-post liability is built on waiting for an adverse event to occur, then assigning liability to the injurer [54]. The legal interpretation of liability already appeared in Roman Law, but the modern framework of product liability first emerged in the early twentieth century [55] in relation to a faulty automobile wheel causing injuries to the driver. The ruling cited the reasonable basis for the automobile manufacturer to know product risks and to make an extra effort to ensure the safety of anyone coming into contact with the car, even if the wheel was produced by another company. The next milestone case in 1932 [56] established duty of care with respect to a person who consumed spoiled ginger beer produced by a bottler and contracted severe gastroenteritis. The bottler had to pay restitution owing to his negligence. Finally, standard of care was put forward in 1947  [57] in a complicated multi-party liability case among the US government, a railroad company, and two boating companies, where the mooring lines of a barge carrying flour belonging to the government were disconnected by another barge, causing damage to other ships and eventually sinking. Liability was shared among the companies; specifically, the boating company operating the barge that cut the lines failed to look after the safety and security of the other barges affected. Interestingly, the judge based his ruling on inadequate risk management: the burden of precautions taken by the company was smaller than the likelihood, and the impact of the incident multiplied together.
@@
Although there were several liability cases against software companies in the last decades, companies often escaped fines using nuances of contract law and limitations of liability and disclaimers in end-user license agreements (EULA) [20]. In fact, the flexible interpretation of acceptable risk taken by the user of a software product made this a gray area. This has changed dramatically towards the end of the last decade with the emergence of the European General Data Protection Regulation (GDPR) and its penalty system for organizations mismanaging the personal data of their users. The GDPR has defined both the duty and standard of care, data protection agencies have collected billions of Euros in fines, and the cost-benefit assessment of cybersecurity-related investments has changed for the better, at least in terms of the safekeeping of personal data. Concurrently, governments worldwide have become increasingly worried about the changing landscape of cyberattacks, which has shown a shift from accessing/exposing user data towards damaging and controlling critical infrastructure integrated across governmental institutions and private organizations. The 2017 NotPetya ransomware [58] and the 2021 SolarWinds attacks [59] stand out as having far-reaching implications for national security, letting malicious attackers breach and potentially take over control of systems of critical importance (governmental organizations, financial institutes, logistics companies, etc.) in the Ukraine, in the United States, and across the globe. These incidents made governments prioritize the cyber-hardening of their national infrastructure and, specifically, focus on software supply chain security. Accordingly, both the United States (via the 2021 Executive Order on Improving the Nation’s Cybersecurity [22]) and the European Union (via the 2022 Cyber Resilience Act [60]) acted swiftly and put forward regulations mandating secure software development, standardized product cybersecurity compliance, and a secure software supply chain [61], and also reducing the information asymmetry pertaining to non-expert end-users of software products. Software liability also factors in the end-user’s technical (lacking expertise to assess product security) and legal inability (EULA prohibiting the inspection of software code) to evaluate software products they use. In fact, the EU has enacted a new Product Liability Directive [18] that came into force on 9th December 2024, with the intent to explicitly designate software as a “product”; this, of course, comes with strict liability clauses. The recent CrowdStrike incident has shown the world that this just might be an effective approach for managing such scenarios (see Section 4 for a detailed case study).
@@
Another intriguing aspect of software liability comes from cyber-insurance. The devastating cybersecurity incidents and the ongoing multi-modal armed conflicts affecting critical infrastructure raised alarms in the insurance industry regarding the insurability of cyber risks [62], potentially inhibiting the market that experienced real growth in recent years after decades of underdevelopment. In fact, insurers, reinsurers, academics, and cybersecurity experts are trying to establish baseline scenarios in the industrial control system domain [63]. Indeed, if the cyber risk is systemic and catastrophic losses are expected in the critical infrastructure sector, the insurance industry will lose interest in underwriting new policies. In that case, governments would shoulder the responsibility to be insurers of last resort, similar to instances of frequent natural disasters [64]. Supposing this happens, the increased scrutiny for software product security laid down in recent regulations is even more sensible [65]. Even so, a potential insurer role would also be a financial and administrative burden for the government, especially taking into account tangled software supply chains. On the other hand, supply chain security solutions and best practices are available; if a company decides to invest money and effort, a much-improved level of security is within reach. Recognizing this, the current United States National Cybersecurity Strategy [66], released in April 2023, proposed a liability waiver mechanism designed to reward software companies willing to undergo and pass a government-mandated product security audit. (Such audits have been studied extensively; we refer the reader to [67] for an overview.) Such a waiver mechanism could just be the missing financial incentive for software companies to change their usual ways regarding secure products.